# Fractionalized superconductivity from Majorana glue in the Kitaev–Kondo lattice

### v1

@@ Acknowledgements.
We would like to thank L. Klebl and J. Profe for their help with the use of their FRG codebase [29] and fruitful physics discussions, as well as J. Knolle for providing the data of Ref. 27. MB would like to thank A. Savvinos for her insights around the group theoretic aspects of this project. This work was funded by the Deutsche Forschungsgemeinschaft (DFG) through SFB 1143 (project id 247310070) and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project id 390858490). S.R. acknowledges support from the Australian Research Council through Grant No. DP200101118 and DP240100168. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project “k102de–FRG”. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the DFG – 440719683.

### v2

@@ Acknowledgements.
We would like to thank L. Klebl and J. Profe for their help with the use of their FRG codebase [29] and fruitful physics discussions, as well as J. Knolle for providing the data of Ref. 27. M.B. would like to thank A. Savvinos for her insights around the group theoretic aspects of this project. This work was funded by the Deutsche Forschungsgemeinschaft (DFG) through SFB 1143 (project id 247310070), SFB 1238 (project id 277146847), the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, project id 390858490) and the Emmy Noether Program (SE 3196/2-1, project id 544397233). S.R. acknowledges support from the Australian Research Council through Grant No. DP200101118 and DP240100168. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project “k102de–FRG”. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the DFG – 440719683.